In q-5 of ,NCERT-12,ex 10.2 "Find λ and μ if      (2i+6j+27k)x(i+λj+μk)=0"
Since the vector product is zero, the two vectors are parallel right? When vectors are parallel, their direction ratios are equal right? So is λ=6 and μ=27?  

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
if vectors are parallel then their direction ratio are in constant proportion.
 
Hence 2/1 = 6/lambda = 27/Mu
 
lambda = 3
Mu = 27/2
Vikas TU
4 years ago
Dear student 
Please follow this link for better understanding 
Good Luck